# Time-Series at the Edge: Tiny Separable CNNs for Wearable Gait Detection and Optimal Sensor Placement

### v1

@@ Abstract
We study on-device time-series analysis for gait detection in Parkinson’s disease (PD) from short windows of triaxial acceleration, targeting resource-constrained wearables and edge nodes. We compare magnitude thresholding to three 1D CNNs for time-series analysis: a literature baseline (separable convolutions) and two ultra-light models—one purely separable and one with residual connections. Using the BioStampRC21 dataset, 2 s windows at 30 Hz, and subject-independent leave-one-subject-out (LOSO) validation on 16 PwPD with chest-worn IMUs, our residual separable model (Model 2, 533 params) attains PR–AUC = 94.5%, $F_{1}=91.2\%$, MCC = 89.4%, matching or surpassing the baseline (5,552 params; PR–AUC = 93.7%, $F_{1}=90.5\%$, MCC = 88.5%) with ${\sim}10{\times}$ fewer parameters. The smallest model (Model 1, 305 params) reaches PR–AUC = 94.0%, $F_{1}=91.0\%$, MCC = 89.1%. Thresholding obtains high recall (89.0%) but low precision (76.5%), yielding many false positives and high inter-subject variance. Sensor-position analysis (train-on-all) shows chest and thighs are most reliable; forearms degrade precision/recall due to non-gait arm motion; naive fusion of all sites does not outperform the best single site. Both compact CNNs execute within tight memory/latency budgets on STM32-class MCUs (sub-10 ms on low-power boards), enabling on-sensor gating of transmission/storage. Overall, ultra-light separable CNNs provide a superior accuracy–efficiency–generalization trade-off to fixed thresholds for wearable PD gait detection and underscore the value of tailored time-series models for edge deployment.
@@ Introduction
Parkinson’s Disease (PD) is now one of the fastest-growing neuro-degenerative disorders in the world (chaudhuri2024economic). Despite optimal treatment, the progression of the disease leads to a gradual deterioration in motor and cognitive disorders, with inevitable compromise in patients’ quality of life. Several works in the literature study gait disabilities using wearable, small, and non-invasive sensors. To design personalized interventions and overcome the limitations related to self-reported questionnaires and diaries, the focus is to improve clinical and home care through the additional information on daily life performance provided by Inertial Measurement Units (IMUs) (corra2021comparison). It is essential to develop algorithms for rapid and precise detection of motor activity in subjects with Parkinson’s Disease to intervene promptly in the detection of anomalies to safeguard the subject’s health. To carry out continuous and constant monitoring of PD subjects, particularly in a domestic and unsupervised environment, the first fundamental step is the detection of the activities carried out by the subject (muthukrishnan2020wearable; corra2021comparison). In particular, gait activities are of interest due to the fact that PD severely affects the gait cycle and, therefore, gait cycle impairments should be monitored and detected promptly. Following gait detection, it is advisable to perform gait analysis to detect all possible anomalies such as fluctuations, dyskinesias, etc (sigcha2020deep; mancini2021measuring). The main studies either process the signal offline or use Cloud-based services. Cloud-based architecture needs data to be moved to the data centers, and are typically employed to support data long-term analyses. However, this scenario is contrary to IoT application requirements, such as limit costs, memory footprint, processing, and communication resources (kianoush2023random). In recent years, there has been a shift towards edge computing, which means processing data on devices closer to the data sources, rather than relying on cloud-based services to reduce the amount of data to be sent to the cloud by the IoT devices (fathalla2022lstm). This shift has also been enabled by advances in embedded systems, as well as tools and accelerators that make it possible to run machine learning and artificial intelligence algorithms on low-power devices (dini2024overview).
@@
In this paper, an in-depth investigation of gait recognition techniques is conducted as a first step towards developing a decentralized architecture for Parkinson’s disease monitoring where distributed processing happens across wearable devices, Edge nodes, and the Cloud. The proposed gait recognition algorithm aims to reduce data transmission through local data processing, optimizing the system’s energy efficiency and improving the quality of stored data by eliminating redundant or irrelevant data samples. Compared to previous works, we provide a comprehensive evaluation that compares simple threshold-based methods to separable CNNs based frameworks, including a Residual Network structure. Our evaluation covers multiple sensor positions and combinations, assesses the methods’ performance across different patients (Leave-One-Subject-Out validation), and examines computational complexity and inference time on a target MCU.
@@
    We introduce two lightweight configurations of 1D-CNNs based architectures achieving a better complexity-performance tradeoff and generalization via F1-score compared to a reference, larger baseline. Through the LOSO evaluation, we show that is possible to design a super-light architecture that matches performances of models 18x times biggers while having extremely light inference times in tasks of binary gait detection on micro-controllers.
@@
    We discuss flaws of a threshold based approach, mainly concerning an inability for the framework to generalize among different subjects and sensor positions, and to maintain a stable accuracy level.
@@ Methods
The raw data utilized in this study was obtained from the ”BioStampRC21” dataset (hereafter, BioStamp) by (PD-BIOSTAMP-RC21), which is accompanied by clinical metadata including patient IDs and health statuses (Patients with Parkinson’s Disease (PwPD) or Control). BioStamp gathers data from 17 PwPD and 17 control subjects of matching ages over an average time of 45.4 hours. The PwPD sample consists of 7 women and 10 men with a mean age of 66.4 years (11.3 sd), while the age-matched sample consists of 13 women and 4 men with a mean age of 64.0 years (9.9 sd). Acceleration data in each individual subject is recorded from 5 different sensor position: chest (ch), left thigh (ll), right thigh (rl), left forearm (lh), right forearm (rh). A clinical annotation file for each subject provides a description of the clinical test being conducted at each timestamp during data collection.
@@
Data annotations were categorized into ”gait” and ”non-gait” events based on the nature of the clinical test. For example, gait-related activities included UPDRS and UHDRS test versions of ”Timed Up and Go”, ”10 Meter Walk Test”, ”Gait”, and ”Tandem Walking”. Events that did not comprise the selected gait activities were categorized as non-gait. The only annotation excluded from both criteria was the ”UHDRS 15 - Retropulsion Pull Test”, due to ambiguity in classifying the event and a negligible relevance both in the dataset and for the representation of domestic activities in PwPD. Sensor data were sourced from accelerometers placed at the five previously described body positions: ch, ll, rl, lh, rh. Each participant’s sensor data underwent column standardization, renaming original columns to maintain consistency across the dataset. Subsequently, the accelerometer readings were uniformly resampled from 31.25 Hz to 30 Hz to allow the use of integer-length windows.
@@
Processed data from each sensor were compiled into structured dictionaries, organized by participant IDs and sensor positions. This led to the creation of 5 datasets, based on sensor positions. During the dataset generation process, subject ”ID38” was excluded from the dataset generation process due to data quality issues. Acceleration data from all 5 sensors were aggregated across each participant and concatenated to form a comprehensive dataset. Each dataset (the 5 per-sensor datasets and the comprehensive dataset) was divided into training set (60%), a validation set (20%), and test set (20%) using stratified random splitting to maintain a consistent label distribution. The stratification ensured balanced representation of gait and non-gait labels in both training and validation datasets.
@@
. To prepare the dataset for LOSO cross-validation, only PwPD were selected for the analysis, resulting in a final pool of 16 subjects: IDs = [6, 10, 12, 13, 15, 17, 23, 24, 25, 33, 35, 36, 40, 42, 44, 63]. For each of the resulting PwPD, data from that participant served as the test set, while the remaining participants data constituted training (70%) and validation (30%) dataset. Accelerometer data only from the chest sensor were extracted for this analysis as they proved to be the most stable from our sensor-dependent analysis. Every datasets obtained – LOSO, sensor-specific and combined – is composed by 60 x 3 shape windows, representing the 2 second windows with triaxial accelerometer data sampled at 30hz.
@@
where $n$ is the number of samples in the window, and $a_{x_{i}}$, $a_{y_{i}}$, and $a_{z_{i}}$ are the measured acceleration components along the 3 measurement axes for the $i^{th}$ time step. The identification of gait windows is performed by evaluating if the magnitude of the window is above the threshold. Consistently to previous works Borzì-2023, the optimal threshold is selected as the value that maximizes the F1 score across the training dataset. In the evaluation process, all windows with magnitude higher than the optimal threshold are classified as gait, and conversely the windows with a lower magnitude are classified as non-gait.
@@
In this section we describe and evaluate three different one-dimensional Convolutional Neural Networks (1D-CNNs) architectures for Time Series Analysis.
@@
Parameter profile. Params (including BN $\gamma,\beta$): SepConv1: $39$; BN1: $16$; SepConv2: $184$; BN2: $32$; Dense: $34$; total $\mathbf{305}$.
@@
Model 2: SuperLight Residual Separable CNN.
The second architecture we analyse, Model 2, keeps the same lightweight design footprint as Model 1 (separable 1D convolutions and two stages) but changes two aspects: (i) each stage is a residual block with an identity skip (projected with a $1{\times}1$ convolution when channel dimensions differ), and (ii) both convolutions use a larger kernel (k=9k{=}9). The first block downsamples with average pooling (factor $2$); the second does not downsample. Each block has convolutions with bias disabled, followed by BatchNorm and ReLU; the residual sum is followed by a ReLU. A global average pooling layer and a two-class softmax head complete the model.
Parameter profile. With input $(60,3)$ the trainable parameters are: Block 1 (SepConv+BN+projection) =107=107, Block 2 =392=392, Dense head =34=34, totaling $\mathbf{533}$ parameters.
@@
All models share the same pipeline and training hyperparameters. Due to compatibility requiremenrs all models have Softmax heads and use categorical cross-entropy and one-hot labels.
Data split (LOSO).
@@
Notation.
Let $\mathrm{TP},\mathrm{TN},\mathrm{FP},\mathrm{FN}$ denote the confusion counts for the positive class (gait), then we compute the following. PR-AUC (Average Precision),i.e., the area under the precision–recall curve (computed as Average Precision); threshold-free and robust to class imbalance. F1 at fixed threshold, i.e., the harmonic mean of precision and recall at the model’s fixed decision threshold tautau, selected on validation by maximizing F1:
@@
Furthermore, in the summary table 4 we include the Geometric Mean of TPR and TNR (GM), the chance-corrected agreement (Cohen’s K), the Accuracy (kept for reference but de-emphasized under class imbalance), and the MCC Matthews correlation coefficient, which provides a single summary that captures whether a model balances both classes rather than succeeding by majority-class guessing.
@@ Results
Table 4: Average LOSO performance. Metrics are percentages; higher is better.
Metric Model 1 Model 2 Baseline Threshold
@@
Table 10 shows the result on the evaluation of different sensor positions, specifically after training the models on the entire dataset and then testing each position separately. Chest and thighs are consistently strongest across models. For the residual separable CNN (Model 2), left-thigh is best (F1 = 88.7%), followed by chest (F1 = 87.5%) and right-thigh (F1 = 84.7%). The lightweight separable CNN (Model 1) follows the same pattern (left-thigh 88.5%, chest 84.2%). The literature baseline also peaks on the left-thigh (89.1%) and chest (88.4%). Forearms underperform for all CNNs (Model 2: 76.0–81.8% F1), and the threshold baseline shows the highest precision drop on forearms (PPV = 43.0%/42.2%; F1 = 57.1%/57.0% for left and right, respectively) despite high recall. The Threshold method performs significantly worse when tested on the fused ’All’ sensor data (F1 = 67.7%) compared to the CNN models (Model 2 F1 = 83.8%), indicating its poor generalization across sensor sites.
Table 5: LOSO evaluation for Model 1 (lightweight SepCNN). Metrics are percentages. $\mu$ and $\sigma$ denote mean and standard deviation across subjects.
@@
Table 6: LOSO evaluation for Model 2 (ResNet). Metrics are percentages. $\mu$ and $\sigma$ denote mean and standard deviation across subjects.
@@
Table 7: LOSO evaluation for the Baseline model. Metrics are percentages. $\mu$ and $\sigma$ denote mean and standard deviation across subjects.
@@
Table 8: LOSO evaluation for the Threshold baseline. Metrics are percentages. $\mu$ and $\sigma$ denote mean and standard deviation across subjects.
@@
We profile three STM32 targets (L4R9I-DISCO, F401RE, H7S78-DK). Table 9 reports memory, MACs, and per-window latency. The Baseline is largest and slowest (133,755 MACs; 13.75 KB RAM; H7: 3.22 ms; L4/F4: 26.68/32.72 ms). Model 1 is the most compact (9,879 MACs; 4.87 KB RAM) and fastest (H7: 0.45 ms; L4/F4: 3.42/4.18 ms). Model 2 offers the best accuracy–efficiency balance (18,303 MACs; H7: 0.70 ms; L4/F4: 5.50/6.80 ms), delivering slightly higher $F_{1}$ than Baseline with $\sim\!7.3\times$ fewer MACs. All models fit without external flash/RAM, enabling on-sensor gating. Because dynamic energy on MCUs scales roughly with operation count, lower MACs imply lower energy per inference (li2025qpart). By contrast, fixed-threshold detection has cost that grows linearly with input length and is effectively negligible on our windows (Borzì-2023), but its poor precision and high inter-subject variance (Table 8) can make downstream false-positive costs (radio, storage, block activation) dominate.
@@
Model Flash RAM MACs Inference Latency (ms) (KB) (KB) (M) L4R9I-DISCO F401RE H7S78-DK Model 1 20.62 4.87 9879 3.42 4.18 0.45 Model 2 23.86 11.93 18303 5.50 6.80 0.70 Baseline 42.12 13.75 133755 26.68 32.72 3.22
Table 10: Train-on-all, per-sensor evaluation (subject-mixed). Metrics are percentages. Acc = accuracy; TNR = specificity; TPR = recall; PPV = precision.

### v2

@@ Abstract
We study on-device time-series analysis for gait detection in Parkinson’s disease (PD) from short windows of triaxial acceleration, targeting resource-constrained wearables and edge nodes. We compare magnitude thresholding to three 1D CNNs for time-series analysis: a literature baseline (separable convolutions) and two ultra-light models—one purely separable and one with residual connections. Using the “BioStampRC21” dataset, 2 s windows at 30 Hz, and subject-independent leave-one-subject-out (LOSO) validation on 16 PwPD with chest-worn IMUs, our residual separable model (Model 2, 533 params) attains PR–AUC = 94.5%, $F_{1}=91.2\%$, MCC = 89.4%, matching or surpassing the baseline (5,552 params; PR–AUC = 93.7%, $F_{1}=90.5\%$, MCC = 88.5%) with ${\sim}10{\times}$ fewer parameters. The smallest model (Model 1, 305 params) reaches PR–AUC = 94.0%, $F_{1}=91.0\%$, MCC = 89.1%. Thresholding obtains high recall (89.0%) but low precision (76.5%), yielding many false positives and high inter-subject variance. Sensor-position analysis (train-on-all) shows chest and thighs are most reliable; forearms degrade precision/recall due to non-gait arm motion; naive fusion of all sites does not outperform the best single site. Both compact CNNs execute within tight memory/latency budgets on STM32-class MCUs (sub-10 ms on low-power boards), enabling on-sensor gating of transmission/storage. Overall, ultra-light separable CNNs provide a superior accuracy–efficiency–generalization trade-off to fixed thresholds for wearable PD gait detection and underscore the value of tailored time-series models for edge deployment.
@@ Introduction
Parkinson’s disease (PD) is now one of the fastest-growing neurodegenerative disorders in the world (chaudhuri2024economic). Despite optimal treatment, the progression of the disease leads to a gradual deterioration in motor and cognitive disorders, with inevitable compromise in patients’ quality of life. Several works in the literature have studied gait disabilities using wearable, small, and non-invasive sensors. To design personalized interventions and overcome the limitations related to self-reported questionnaires and diaries, the focus is to improve clinical and home care through the additional information on daily life performance provided by Inertial Measurement Units (IMUs) (corra2021comparison). It is essential to develop algorithms for rapid and precise detection of motor activity in subjects with Parkinson’s disease to intervene promptly in the detection of anomalies to safeguard the subject’s health. To carry out continuous and constant monitoring of PD subjects, particularly in a domestic and unsupervised environment, the first fundamental step is the detection of the activities carried out by the subject (muthukrishnan2020wearable; corra2021comparison). In particular, gait activities are of interest due to the fact that PD severely affects the gait cycle and, therefore, gait cycle impairments should be monitored and detected promptly. Following gait detection, it is advisable to perform gait analysis to detect all possible anomalies such as fluctuations, dyskinesias, etc. (sigcha2020deep; mancini2021measuring). The main studies either process the signal offline or use Cloud-based services. Cloud-based architectures need data to be moved to data centers, and are typically employed to support long-term data analyses. However, this scenario is contrary to IoT application requirements, such as limiting costs, memory footprint, processing, and communication resources (kianoush2023random). In recent years, there has been a shift towards edge computing, which means processing data on devices closer to the data sources, rather than relying on cloud-based services to reduce the amount of data to be sent to the cloud by the IoT devices (fathalla2022lstm). This shift has also been enabled by advances in embedded systems, as well as tools and accelerators that make it possible to run machine learning and artificial intelligence algorithms on low-power devices (dini2024overview).
@@
In this paper, an in-depth investigation of gait recognition techniques is conducted as a first step towards developing a decentralized architecture for Parkinson’s disease monitoring where distributed processing happens across wearable devices, edge nodes, and the cloud. The proposed gait recognition algorithm aims to reduce data transmission through local data processing, optimizing the system’s energy efficiency and improving the quality of stored data by eliminating redundant or irrelevant data samples. Compared to previous works, we provide a comprehensive evaluation that compares simple threshold-based methods to separable CNNs-based frameworks, including a Residual Network structure. Our evaluation covers multiple sensor positions and combinations, assesses the methods’ performance across different patients (Leave-One-Subject-Out validation), and examines computational complexity and inference time on a target MCU.
@@
    We introduce two lightweight configurations of 1D-CNNs-based architectures achieving a better complexity-performance tradeoff and generalization via F1-score compared to a reference, larger baseline. Through the LOSO evaluation, we show that it is possible to design a super-light architecture that matches the performance of models 18× larger while having extremely light inference times in tasks of binary gait detection on microcontrollers.
@@
    We discuss flaws of a threshold-based approach, mainly concerning an inability for the framework to generalize among different subjects and sensor positions, and to maintain a stable accuracy level.
@@ Methods
The raw data utilized in this study was obtained from the “BioStampRC21” dataset (hereafter, BioStamp) by (PD-BIOSTAMP-RC21), which is accompanied by clinical metadata including patient IDs and health statuses (patients with Parkinson’s disease (PwPD) or Control). BioStamp gathers data from 17 PwPD and 17 control subjects of matching ages over an average time of 45.4 hours. The PwPD sample consists of 7 women and 10 men with a mean age of 66.4 years (11.3 SD), while the age-matched sample consists of 13 women and 4 men with a mean age of 64.0 years (9.9 SD). Acceleration data in each individual subject is recorded from 5 different sensor positions: chest (ch), left thigh (ll), right thigh (rl), left forearm (lh), right forearm (rh). A clinical annotation file for each subject provides a description of the clinical test being conducted at each timestamp during data collection.
@@
Data annotations were categorized into “gait” and “non-gait” events based on the nature of the clinical test. For example, gait-related activities included UPDRS and UHDRS test versions of “Timed Up and Go”, “10 Meter Walk Test”, “Gait”, and “Tandem Walking”. Events that did not comprise the selected gait activities were categorized as non-gait. The only annotation excluded from both criteria was the “UHDRS 15 - Retropulsion Pull Test”, due to ambiguity in classifying the event and a negligible relevance both in the dataset and for the representation of domestic activities in PwPD. Sensor data were sourced from accelerometers placed at the five previously described body positions: ch, ll, rl, lh, rh. Each participant’s sensor data underwent column standardization, renaming original columns to maintain consistency across the dataset. Subsequently, the accelerometer readings were uniformly resampled from 31.25 Hz to 30 Hz to allow the use of integer-length windows.
@@
Processed data from each sensor were compiled into structured dictionaries, organized by participant IDs and sensor positions. This led to the creation of 5 datasets, based on sensor positions. Subject “ID38” was excluded during the dataset generation process due to data quality issues. Acceleration data from all 5 sensors were aggregated across each participant and concatenated to form a comprehensive dataset. Each dataset (the 5 per-sensor datasets and the comprehensive dataset) was divided into training set (60%), a validation set (20%), and test set (20%) using stratified random splitting to maintain a consistent label distribution. The stratification ensured balanced representation of gait and non-gait labels in both training and validation datasets.
@@
To prepare the dataset for LOSO cross-validation, only PwPD were selected for the analysis, resulting in a final pool of 16 subjects: IDs = [6, 10, 12, 13, 15, 17, 23, 24, 25, 33, 35, 36, 40, 42, 44, 63]. For each of the resulting 16 PwPD, data from that participant served as the test set, while the remaining participants’ data constituted training (70%) and validation (30%) datasets. Accelerometer data only from the chest sensor were extracted for this analysis as they proved to be the most stable from our sensor-dependent analysis. Every dataset obtained – LOSO, sensor-specific and combined – is composed of windows of shape $60\times 3$, representing the 2-second windows with triaxial accelerometer data sampled at 30 Hz.
@@
where $n$ is the number of samples in the window, and $a_{x_{i}}$, $a_{y_{i}}$, and $a_{z_{i}}$ are the measured acceleration components along the 3 measurement axes for the $i^{th}$ time step. The identification of gait windows is performed by evaluating whether the magnitude of the window is above the threshold. Consistent with previous works Borzì-2023, the optimal threshold is selected as the value that maximizes the F1 score across the training dataset. In the evaluation process, all windows with magnitude higher than the optimal threshold are classified as gait, and conversely the windows with a lower magnitude are classified as non-gait.
@@
In this section we describe and evaluate three different one-dimensional Convolutional Neural Network (1D-CNNs) architectures for time-series analysis.
@@
Parameter profile Params (including BN $\gamma,\beta$): SepConv1: $39$; BN1: $16$; SepConv2: $184$; BN2: $32$; Dense: $34$; total $\mathbf{305}$.
@@
Model 2: Super-Light Residual Separable CNN
The second architecture we analyze, Model 2, keeps the same lightweight design footprint as Model 1 (separable 1D convolutions and two stages) but changes two aspects: (i) each stage is a residual block with an identity skip (projected with a $1{\times}1$ convolution when channel dimensions differ), and (ii) both convolutions use a larger kernel (k=9k{=}9). The first block downsamples with average pooling (factor $2$); the second does not downsample. Each block has convolutions with bias disabled, followed by BatchNorm and ReLU; the residual sum is followed by a ReLU. A global average pooling layer and a two-class softmax head complete the model.
Parameter profile With input $(60,3)$ the trainable parameters are: Block 1 (SepConv+BN+projection) =107=107, Block 2 =392=392, Dense head =34=34, totaling $\mathbf{533}$ parameters.
@@
All models share the same pipeline and training hyperparameters. Due to compatibility requirements all models have Softmax heads and use categorical cross-entropy and one-hot labels.
Data split (LOSO)
@@
Notation
Let $\mathrm{TP},\mathrm{TN},\mathrm{FP},\mathrm{FN}$ denote the confusion counts for the positive class (gait), then we compute the following. PR-AUC (Average Precision), i.e. the area under the precision–recall curve (computed as Average Precision); threshold-free and robust to class imbalance. F1 at fixed threshold, i.e., the harmonic mean of precision and recall at the model’s fixed decision threshold $\tau$, selected on validation by maximizing F1:
@@
Furthermore, in the summary table 4 we include the Geometric Mean of TPR and TNR (GM), the chance-corrected agreement (Cohen’s K), the Accuracy (kept for reference but de-emphasized under class imbalance), and the Matthews correlation coefficient (MCC), which provides a single summary that captures whether a model balances both classes rather than succeeding by majority-class guessing.
@@ Results
Table 4: Average LOSO performance.
Metric % Model 1 Model 2 Baseline Threshold
@@
Table 10 reports per-sensor performance, with all models trained on the full dataset and evaluated on each position. Chest and thighs are strongest across models. For the residual separable CNN (Model 2), left-thigh is best (F1 = 88.7%), followed by chest (F1 = 87.5%) and right-thigh (F1 = 84.7%). The lightweight separable CNN (Model 1) follows the same pattern (left-thigh 88.5%, chest 84.2%). The literature baseline also peaks on the left-thigh (89.1%) and chest (88.4%). Forearms underperform for all CNNs (Model 2: 76.0–81.8% F1), and the threshold baseline shows the highest precision drop on forearms (PPV = 43.0%/42.2%; F1 = 57.1%/57.0% for left and right) despite high recall. The Threshold method performs markedly worse when tested on the fused ’All’ sensor data (F1 = 67.7%) compared to the CNN models (Model 2 F1 = 83.8%), indicating its poor generalization across sensors.
Table 5: LOSO evaluation for Model 1.
@@
Table 6: LOSO evaluation for Model 2 (ResNet).
@@
Table 7: LOSO evaluation for the Baseline model.
@@
Table 8: LOSO evaluation for the Threshold baseline.
@@
We profile three STM32 targets (L4R9I-DISCO, F401RE, H7S78-DK). Table 9 reports memory, MACs, and per-window latency. The Baseline is largest and slowest (133,755 MACs; 13.75 KB RAM; H7: 3.22 ms; L4/F4: 26.68/32.72 ms). Model 1 is the most compact (9,879 MACs; 4.87 KB RAM) and fastest (H7: 0.45 ms; L4/F4: 3.42/4.18 ms). Model 2 offers the best accuracy–efficiency balance (18,303 MACs; H7: 0.70 ms; L4/F4: 5.50/6.80 ms), delivering slightly higher $F_{1}$ than Baseline with $\sim\!7.3\times$ fewer MACs. All models fit without external flash/RAM, enabling on-sensor gating. Because dynamic energy on MCUs scales roughly with operation count, lower MACs imply lower energy per inference (li2025qpart). By contrast, fixed-threshold detection has a cost that grows linearly with input length and is effectively negligible on our windows (Borzì-2023), but its poor precision and high inter-subject variance (Table 8) can make downstream false-positive costs (radio, storage, block activation) dominate.
@@
Model Flash RAM MACs Inference Latency (ms) (KB) (KB) L4R9I-DISCO F401RE H7S78-DK Model 1 20.62 4.87 9879 3.42 4.18 0.45 Model 2 23.86 11.93 18303 5.50 6.80 0.70 Baseline 42.12 13.75 133755 26.68 32.72 3.22
Table 10: Train-on-all, per-sensor evaluation (subject-mixed). TNR = specificity; TPR = recall; PPV = precision.